OpenStudy (anonymous):

Substitution x= -1+2y and x=9-2y

OpenStudy (anonymous):

is x=4.5 and y=11/d the answer ?

OpenStudy (hockeychick23):

no, x=4 and y= 2.5, if I were you I would just set the equations equal to each other for this, its a lot easier

OpenStudy (anonymous):

do you round your answer ?

OpenStudy (hockeychick23):

Nope, thats the final answer you can write it as -1+2y = 9- 2y Then you should get all the y's on on side and all of the numbers on the other so you can subtract 2y from both sides